OpenStudy (cedric):

Solve the inequality 2(x + 6) 3x - 2 and show all of your work. You can use >= to represent the greater than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

OpenStudy (anonymous):

I think you need an inequality sign between 2(x+6) and 3x-2

OpenStudy (anonymous):

I need it to help you solve the problem

OpenStudy (cedric):

ok hold on ill get it

OpenStudy (cedric):

>=

OpenStudy (anonymous):

Okay. So first you need to simplify the left side of the equation. So multiply 2 with (x+6) by distributing inside the parentheses: 2x+12>=3x-2 Now in order to solve the inequality you need to isolate the variable, can you do that?

OpenStudy (cedric):

uhm ok you'd do that by factoring the two to both sides right?

OpenStudy (anonymous):

Nope, you just subtract the 2x from both sides and add the two to both sides

OpenStudy (cedric):

so now theres x + 10 >= 1x

OpenStudy (cedric):

oh forgot to aaddd

OpenStudy (anonymous):

Nope. 2x + 12 >= 3x - 2 12>=x - 2 x<=14 because the 2 is negative you have to add and not subtract it to cancel it from the right side, and subtracting the 2x cancels it from the left side

OpenStudy (cedric):

oh you meant both sides of that one problem ohhhh